OpenStudy (anonymous):

please help. Change this radical to an algebraic expression with fractional exponents. 5√x^3

OpenStudy (anonymous):

\[\huge \sqrt[m]{x^n}=x^{\frac{n}{m}}\]

OpenStudy (anonymous):

so is it x(3/5)?